Please sign and return with the exam:

I, , affirm that I will not discuss any aspect of this exam including, but not limited to:

• its form

• its content

• its degree of difficulty

• the amount of time you took to complete it

with anyone except Prof. Little, until after 12:00 noon on Friday, April 21.